OpenStudy (anonymous):

need the step by step please. x^2-3x-1=0

OpenStudy (anonymous):

use the quadratic formula

jhonyy9 (jhonyy9):

x*2 -3x -1 =0 3 +/- radical ( (-3)*2 - 4(1)(-1) ) 3 +/- radical (9+4) x_1,_2 =-------------------------- = ---------------- = 2(1) 2 3 +/- radical 13 = --------------- 2

OpenStudy (anonymous):

\[-b+or-\sqrt{b^2-4ac}\] --------------------- \[2a\]

OpenStudy (anonymous):

she said step by step....lol

OpenStudy (anonymous):

well you have to try it atleast you have the formula

OpenStudy (anonymous):

use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] put \[a=1,b=-3,c=-1\] get \[\frac{-(-3)\pm\sqrt{(-3)^2-4\times 1\times -1}}{2\times 1}\]

OpenStudy (anonymous):

this gives \[\frac{9\pm\sqrt{9+4}}{2}\]

OpenStudy (anonymous):

which in turn gives \[\frac{9\pm\sqrt{13}}{2}\] and that is your answer

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

That's much better, lol

jhonyy9 (jhonyy9):

hey satellite73 thos are not correct solutions

jhonyy9 (jhonyy9):

please check you newly

OpenStudy (anonymous):

jhonny is right

OpenStudy (anonymous):

at least he presented step by step, but lol

jhonyy9 (jhonyy9):

ty heromiles

OpenStudy (anonymous):

I hope levie isn't confused by any of this...lol